OpenStudy (anonymous):

A rectangular shaped living room has a length of 2x+6. If the area of the living room is 1/2, what is the width of the living room? A) 4x+12 B) -2x-5.5 C) x+3 D) 1/2x + 6 or E) 1/4x+12 ?

myininaya (myininaya):

\[LW=A\] \[(2x+6)W=\frac{1}{2}\]

myininaya (myininaya):

\[\frac{1}{2x+6}*(2x+6)W=\frac{1}{2}*\frac{1}{2x+6}\] \[W=\frac{1}{2(2x+6)}=\frac{1}{4x+12}\]

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

Oh i get it thank you!

myininaya (myininaya):

np